If logπ‘Ž π‘ž = 1/5 solve the equation , π‘₯ logπ‘Ž π‘ž + (π‘₯ + 2) logπ‘Ž π‘ž^2 = 3

1
Expert's answer
2021-11-09T01:09:04-0500

"xlog_aq+(x+2)og_aq^2=3\\x\times \frac{1}{5}+2(x+2)log_aq=3\\x\times\frac{1}{5}+2(x+2)\times\frac{1}{5}=3\\\frac{1}{5}(x+2x+40=3\\3x+4=15\\x=\frac{11}{3}"